The D string of a guitar is stretched to a tension of 150 N. The string has a mass of 3.4 g and is stretched to a length of 0.69 m. What is the speed of waves on the string?


1
Expert's answer
2020-07-06T15:39:43-0400
"v=\sqrt{\frac{Tl}{m}}\\v=\sqrt{\frac{150\cdot0.69}{0.0034}}=174\frac{m}{s}"